# Exploring Holomorphic Retracts

### v6

@@ 2.4. One dimensional Retracts
Hence we get that $|L(z)|\leq h(z)$ for all $z\in\mathbb{C}^{N}$. In-particular, $|L(a)|\leq h(a)$. But then the very definition of $L$ shows that $|L(a)|\geq h(a)$ and we thereby get that $|L(a)|=h(a)$ (and $ia$ lies in the kernel of $f$ though this observation is not necessary for the proof per se). Hence we have proved that $D$ is convex at $a/h(a)$ according to definition 2.30. ∎
@@
Let $D$ be a balanced pseudoconvex domain in $\mathbb{C}^{N}$ and $a\in D$ such that $D$ is not convex at $a/h(a)$. Then there does not exist any one-dimensional retract of $D$ passing through 00 and $a$.
@@
Assume the contrary, there exists a one-dimensional retract $R$ of $D$ passing through 00 and $a$. By theorem 2.29, $R$ is image of some complex C-geodesic $\varphi:\Delta\to D$. But by lemma 11.1.2 in [Jrncki_invrnt_dst] (see also remark 11.1.3(b) therein), there is no complex $C$-geodesic in $D$. This leads to a contradiction. ∎
Proposition 2.34.
Let $D$ be a balanced pseudoconvex domain in $\mathbb{C}^{N}$ and $a\in D$ such that $D$ is convex at $a/h(a)$. Then $\varphi:\Delta\to D$ defined by $\varphi(\lambda)=\lambda a/h(a)$ is a complex $C$-geodesic on $D$. In particular, there always exists a one dimensional linear retract of $D$ passing through 00 and $a$.
Proof.
Let $\lambda$ be any non-zero element in $\Delta$. Since $D$ is convex at $a/h(a)$ and by proposition 2.3.1(b) in [Jrncki_invrnt_dst], we get
$c_{D}^{*}(\varphi(0),\varphi(\lambda))=c_{D}^{*}(0,\lambda a/h(a))=h(\lambda a/h(a))=|\lambda|$
Note that $c_{D}=\tanh^{-1}c_{D}^{*}$. So
$c_{D}(\varphi(0),\varphi(\lambda))=\tanh^{-1}\left(c_{D}^{*}(\varphi(0),\varphi(\lambda))\right)=\tanh^{-1}(|\lambda|)=p(0,\lambda)$
By proposition 11.1.4 in [Jrncki_invrnt_dst], $\varphi$ is a complex $C$-geodesic. ∎
@@
Combining theorem 2.29 and theorem 3.2 in [Vigue_Cnvx], we obtain the following result:
@@ 2.5. Fixed point sets and Retracts
Let $D$ be a bounded taut domain in $\mathbb{C}^{N}$. Assume that $f:D\to D$ be a holomorphic map which is not an automorphism and $\text{Fix}(f)$ is a non-empty subset of $D$ which consists of more than one point. Then there exists a non-trivial retract $Z$ of $D$ such that ${\rm Fix}(f)\subset Z$
@@
Consider the sequence of holomorphic maps on $D$ given by $f_{n}:=f^{\circ n}$ which denotes as usual the $n$-fold composition of $f$ with itself. Since $\text{Fix}(f)\neq\phi$, $(f_{n})$ is not compactly divergent. By theorem 2.1.29 in [Abate_Itrtn], there exists a retract $Z$ of $D$ and a holomorphic retraction $\rho:D\to Z$ such that $\rho$ is a limit point of the sequence $\{f_{n}\}$. Note that $\rm Fix(f)\subset\rm Z$. $\rm dim(Z)\neq 0$, because $Z$ is a connected complex manifold and #$\#(Fix(f))>1$. By Corollary 2.1.30 in [Abate_Itrtn], $\rm dim(Z)$ is equal to the number of eigenvalues of $df_{z_{0}}$ contained in $\partial\Delta$. If all eigenvalues of $df_{z_{0}}$ are in $\partial\Delta$, then by theorem 2.1.21 (v) in [Abate_Itrtn], we get $f$ is an automorphism. This contradicts the hypothesis. Hence we can conclude that $\rm dim(Z)<N$. This proves that $Z$ is a non-trivial retract of $D$ and $\rm Fix(f)\subset Z$. ∎
3. Proofs of theorems 1.3, 1.4, 1.8, 1.13, 1.15
We would like to admit at the outset that the proof of theorem 1.15 on retracts of polyballs, closely follows the recent work [BBMV-union]; but we give here a thorough treatment even if, at the cost of repetition of essential ideas, for it allows for a better exposition of the proof of both this theorem, as well as perhaps more importantly: proofs of theorems prior to it as in the title of this section, which are fundamental to this article.
Our goal here is to characterize the holomorphic retracts of the Cartesian product $B_{1}\times\ldots\times B_{N}$ (which will be referred to as polyballs for short), where the $B_{j}$’s are bounded balanced strictly convex domains or equivalently, the unit ball with respect to some norm in $\mathbb{C}^{n_{j}}$ for some $n_{j}\in\mathbb{N}$, for each $j=1,\ldots,N$; as noted in the previous section, this is equivalent to their boundaries being $\mathbb{R}$-extremal. As it turns out that the essential arguments for the case of a product of a pair of such balls extend to the more general case, we shall for simplicity of exposition, focus attention for the most part on the case of a product of $2$ such balls. So, let $B_{1}\subset\mathbb{C}^{N_{1}},B_{2}\subset\mathbb{C}^{N_{2}}$ (where $N_{1},N_{2}\in\mathbb{N}$ with say, $\min\{N_{1},N_{2}\}=N_{1}$) denote a pair of bounded balanced convex domains with $\mathbb{R}$-extremal boundaries; in-particular, $B_{1},B_{2}$ are balls with respect to some norms (on $\mathbb{C}^{N_{1}},\mathbb{C}^{N_{2}}$ respectively) and needless to say are domains of holomorphy. The simplest example of such a domain is the standard unit disc in $\mathbb{C}$ which we denote by $\Delta$. However, many nice properties of the disc may well not hold for such domains (i.e., bounded balanced convex domains). To begin with, balls in the $l^{p}$-norm for $1<p<\infty$, while satisfying the condition that their boundary points are $\mathbb{R}$-extremal, need not always be infinitely smooth – they are atmost $[p]$-smooth, where []delimited-[][$\cdot$][ ⋅ ] denotes the greatest integer function. They are thus atleast $C^{1}$-smooth and to get an example where this $C^{1}$-smoothness fails, we may take an intersection of a pair of suitable dilates of the ball in the $l^{p_{1}}$ norm with that of the $l^{p_{2}}$ norm, for a pair of positive distinct reals $p_{1},p_{2}$ in the interval $(1,\infty)$ to get examples of domains whose boundaries are not $C^{1}$-smooth but nevertheless satisfy the $\mathbb{R}$-extremal condition; as this is essentially the well-known fact that the intersection of a pair of strictly convex domains is again strictly convex, we shall provide details with $\mathbb{R}$-extremality replaced by $\mathbb{C}$-extremality, in the following.

### v7

@@ 2.5. One dimensional Retracts
Hence, we get that $|L(z)|\leq h(z)$ for all $z\in\mathbb{C}^{N}$. In-particular, $|L(a)|\leq h(a)$. But then the very definition of $L$ shows that $|L(a)|\geq h(a)$ and we thereby get that $|L(a)|=h(a)$ (and $ia$ lies in the kernel of $f$ though this observation is not necessary for the proof per se). Hence, we have proved that $D$ is convex at $a/h(a)$ according to definition 2.34. ∎
We also note that the notion of holomorphic extremality is equivalent to the notion of $\mathbb{C}$-extremality of the boundary in the presence of convexity, as clarified in the following proposition.
Proposition 2.37.
Let $D$ be a bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$. Assume that $D$ is convex near $p\in\partial D$. Then $D$ is $\mathbb{C}$-extremal at $p$ if and only if $D$ is holomorphic extreme at $p$.
Proof.
Since every holomorphic extreme boundary point of $D$ is $\mathbb{C}$- extremal, it suffices to show that if $D$ is $\mathbb{C}$-extremal at $p$ then $D$ is holomorphic extreme at $p$. Assume that $D$ is $\mathbb{C}$-extremal at $p$. Let $\text{co}(D)$ denote the convex hull of $D$. Firstly note that that $\text{co}(D)$ is a balanced convex domain. Next, we divide the proof into the following steps:
Step-1: If $p\in\partial D$ and $D$ is convex at $p$ then $p\in\partial\text{co}(D)$.
By proposition 2.3.1 (b) in [Jrncki_invrnt_dst], we get that $\gamma_{D}(0;p)=h(p)=1$. Applying proposition 2.3.1 (d), we obtain that $\hat{h}(p)=1$, where $\hat{h}$ is the Minkowski function of $\text{co}(D)$. This shows that $p\in\partial\text{co}(D)$.
Step-2: $p$ is a $\mathbb{C}$-extremal boundary point of $\text{co}(D)$.
To prove this by contradiction, assume that $p$ is not $\mathbb{C}$-extremal boundary point of $\text{co}(D)$. Then there exists non-zero $u\in\mathbb{C}^{N}$ such that $p+tu\in\text{co}(D)$ for all $t\in\Delta$. After applying proposition 2.3.1(a) in [Jrncki_invrnt_dst] we get that $\gamma_{D}(0,p+tu)=1$. Since $D$ is convex near $p$, we may assume that $D$ is convex at $(p+tu)/h(p+tu)$ for all $t\in\Delta$; the equivalence of (iv) and (v) in proposition 2.3.1 in [Jrncki_invrnt_dst] shows that $h(p+tu)=1$ for all $t\in\Delta$. This implies that $p+tu\in\partial D$ for all $t\in\Delta$, which contradicts the fact that $p$ is $\mathbb{C}$-extremal point of $D$. This finishes the proof of Step 2.
Thereafter, applying lemma 11.3.7 in [Jrncki_invrnt_dst] to $\text{co}(D)$, we obtain that $p$ is a holomorphic extreme point of $\text{co}(D)$. Using arguments as in the proof of Step 2, we obtain that $p$ is a holomorphic extreme point of $D$. This finishes the proof of the proposition. ∎
@@
Let $D$ be a balanced pseudoconvex domain in $\mathbb{C}^{N}$ and $a\in D$ such that $D$ is not convex at $a/h(a)$. Then there does not exist any one-dimensional retract of $D$ passing through 0 and $a$.
A proof of this will be provided later in the course of other things in the next section. We next get to a positive result.
Proposition 2.39.
Let $D$ be a balanced pseudoconvex domain in $\mathbb{C}^{N}$ and $a\in D$ such that $D$ is convex at $a/h(a)$. Then there always exists a one dimensional linear retract of $D$ passing through 0 and $a$.
@@
Since $D$ is convex at $p$, there exists a $\mathbb{C}$-linear functional $L:\mathbb{C}^{N}\to\mathbb{C}$ such that $|L(z)|\leq h(z)$ for all $z\in\mathbb{C}^{N}$ and $|L(p)|=1$. By multiplying $L$ by a suitable unimodular constant, we may assume that $L(p)=1$. Consider the linear transformation $T:\mathbb{C}^{N}\to\mathbb{C}^{N}$ defined by $T(z)=L(z)p$. Note that $T$ maps $D$ into $D$ because
$h(T(z))=h(L(z)p)=|L(z)|h(p)\leq h(z)<1.$
Moreover, $T$ is a linear projection map:
T(T(z))=T(L(z)p)=L(L(z)p)p=L(z)L(p)p=L(z)p=T(z).T(T(z))=T(L(z)p)=L(L(z)p)p=L(z)L(p)p=L(z)p=T(z).
This shows that $\text{span}(p)\cap D$ is a linear retract of $D$. Since $D$ is convex at $p$, there exists a $\mathbb{C}$-linear functional $L:\mathbb{C}^{N}\to\mathbb{C}$ such that $|L(z)|\leq h(z)$ for all $z\in\mathbb{C}^{N}$ and $|L(p)|=1$. By multiplying $L$ by a suitable unimodular constant, we may assume that $L(p)=1$. Consider the linear transformation $T:\mathbb{C}^{N}\to\mathbb{C}^{N}$ defined by $T(z)=L(z)p$. Note that $T$ maps $D$ into $D$ because
$h(T(z))=h(L(z)p)=|L(z)|h(p)\leq h(z)<1.$
Moreover, $T$ is a linear projection map:
T(T(z))=T(L(z)p)=L(L(z)p)p=L(z)L(p)p=L(z)p=T(z).T(T(z))=T(L(z)p)=L(L(z)p)p=L(z)L(p)p=L(z)p=T(z).
This shows that $\text{span}(p)\cap D$ is a linear retract of $D$. ∎
Remark 2.40.
  1. (1)
    In the case of bounded balanced convex domains, the geometric Hahn-Banach theorem ensures that every one-dimensional subspace of $D$ is a retract of $D$.
  2. (2)
    Every one dimensional retract $Z$ of $D$ passing through the origin and $a$ can be parametrized by $\lambda a/h(a)$ for $\lambda\in\Delta$.
@@
Combining theorem 2.33 and theorem 3.2 in [Lemp_geodesc], the following result is obtained.
@@
Remark 2.43.
Assume for this remark that $D$ is convex. The existence in abundance of one dimensional retracts of strictly convex domains (in $\mathbb{C}^{N}$) is well-known due to the work of Lempert. Another well-known work which obtains retracts of higher dimensions under certain conditions is [Ian_Grhm] of Graham; as noted therein, the condition in the above theorem that $D_{L}=D\cap L$ be a linear retract of $D$ which is equivalent to the existence of a norm one projection onto $L$ (denoted by $\pi_{L}$ in the proof), is always met if $L$ is one dimensional – this follows by the Hahn–Banach theorem which may be used to first get a linear functional $f$ on $(\mathbb{C}^{N},\mu)$ of norm one which takes the value $1$ at a boundary point of $D_{L}$ (one may view $f$ as mapping $D$ onto $D_{L}\subset L\simeq\mathbb{C}$); the desired projection is then obtained by taking the linear projection of $\mathbb{C}^{N}$ onto $L$ along $\ker(f)$. On the other hand (also noted therein), if there exists a norm one projection of the Banach space $(\mathbb{C}^{N},\mu)$ onto a every linear subspace, then $(\mathbb{C}^{N},\mu)$ is actually a Hilbert space i.e., $\mu$ is actually induced by an inner product owing to classical results mentioned in the introduction namely that of Kakutani in [Kakutni] for the real case and Bohnenblust in [Bhnblst] for the complex case. Thus, not every linear subspace intersected with $D$ need be a retract of $D$.
@@ 2.6. Fixed point sets and Retracts
Let $D$ be a bounded taut domain in $\mathbb{C}^{N}$. Assume that $f:D\to D$ be a holomorphic map which is not an automorphism and $\text{Fix}(f)$ is a non-empty subset of $D$ which consists of more than one point. Then there exists a non-trivial retract $Z$ of $D$ such that ${\rm Fix}(f)\subset Z$.
@@
Consider the sequence of holomorphic maps on $D$ given by $f_{n}:=f^{\circ n}$ which denotes as usual the $n$-fold composition of $f$ with itself. Since $\text{Fix}(f)\neq\phi$, $(f_{n})$ is not compactly divergent. By theorem 2.1.29 in [Abate_Itrtn], there exists a retract $Z$ of $D$ and a holomorphic retraction $\rho:D\to Z$ such that $\rho$ is a limit point of the sequence $\{f_{n}\}$. Note that $\rm Fix(f)\subset\rm Z$. Note also that $\rm dim(Z)\neq 0$, because $Z$ is a connected complex manifold and #(Fix(f))>1\#(Fix(f))>1. By Corollary 2.1.30 in [Abate_Itrtn], $\rm dim(Z)$ is equal to the number of eigenvalues of $df_{z_{0}}$ contained in $\partial\Delta$. If all eigenvalues of $df_{z_{0}}$ are in $\partial\Delta$, then by theorem 2.1.21 (v) in [Abate_Itrtn], we get $f$ is an automorphism. This contradicts the hypothesis. Hence, we can conclude that $\rm dim(Z)<N$. This proves that $Z$ is a non-trivial retract of $D$ and $\rm Fix(f)\subset Z$. ∎
2.7. Hyperconvex Domains
In this subsection, we briefly recall a standard class of pseudoconvex domains containing convex domains, namely hyperconvex domains. Hyperconvexity is an intermediate notion between convexity and pseudoconvexity that is a generalization of a key concept (going by the name of barriers) characterizing solvability of the Dirichlet problem in dimension one. Owing to this, we recall the connection between hyperconvexity and psh (plurisubharmonic) barrier boundary points. To begin with, let us recall the definition of a psh barrier point from [barrier].
Definition 2.45.
A boundary point $a$ of an domain $D\subset\mathbb{C}^{N}$ is said to be a (global) psh barrier point for $D$ if there is a negative plurisubharmonic function $\varphi$ on $D$ with $\lim_{D\ni z\to a}\varphi(z)=0$. We call $a$ a local psh barrier point for $D$ if there is a (connected) open neighborhood $U$ of $a$ such that $a$ is a psh barrier point for $D\cap U$. We shall say that $D$ has psh barrier boundary (resp. local psh barrier boundary) if all of its boundary points are psh (resp. local psh) barrier points.
Remark 2.46.
Examples of domains whose boundary satisfying the local psh barrier condition, as in the definition, abound. All bounded strictly pseudoconvex domains, more generally, all bounded (weakly) pseudoconvex domains of finite type, serve as examples. More interestingly, non-examples are obtained by looking at domains whose boundaries contain one or more piece of analytic varieties. For instance, while all analytic polyhedra $D$ satisfy this condition; as soon as we remove any non-trivial analytic set $A$, $D\setminus A$ no longer satisfies this condition at any of the (boundary) points $a\in A$.
Let us recall the definition of a hyperconvex manifold from [Hyp-cnvx].
Definition 2.47.
A complex manifold $M$ is said to be hyperconvex if there exists a bounded strictly plurisubharmonic (psh) exhaustion function $\varphi$ on $M$.
Remark 2.48.
  • (i)
    If $F:M_{1}\to M_{2}$ is a proper holomorphic map between complex manifolds $M_{1}$ and $M_{2}$ with $M_{2}$ is hyperconvex, then any bounded strictly plurisubharmonic exhaustion function $\varphi$ on $M_{2}$ induces a bounded strictly plurisubharmonic exhaustion function on $M_{1}$ via the pullback $\psi:=\varphi\circ F$. Hence, $M_{1}$ is hyperconvex. In particular, this shows that hyperconvexity is a biholomorphic invariant.
  • (ii)
    Every hyperconvex domain in $\mathbb{C}^{N}$ is pseudoconvex (see theorem 5.2.1 and proposition 5.2.2 in [Kobyshi]).
  • (iii)
    In the case of bounded balanced pseudoconvex domains, hyperconvexity is equivalent to the continuity of their Minkowski functionals (see remark 3.2.3 a), c) in [Jrncki_invrnt_dst]).
@@ 2.7. Hyperconvex Domains
Every retract of a hyperconvex complex manifold is hyperconvex.
Proof.
Let $Z$ be a retract of a hyperconvex complex manifold $X$ and let $\varphi:X\to[-\infty,c)$ be a bounded plurisubharmonic exhaustion on $X$. Note that $\varphi_{|_{Z}}$ serves as a bounded plurisubharmonic exhaustion function on $Z$. Hence, $Z$ is a hyperconvex complex manifold. ∎
The following theorem (see theorem 1.17.5 in [Jrnck_Frst_stps]) describes the relationship between hyperconvexity and the existence of psh barriers.
Theorem 2.50.
Let $D\subset\mathbb{C}^{N}$ be a domain. Then $D$ is hyperconvex if and only if for any boundary point $a\in\partial D$ (including $a=\infty$ when $D$ is unbounded), $a$ is a psh barrier point for $D$.
Remark 2.51.
If $D$ as in the above theorem is bounded, then by theorem 12.4.7 in [Jrncki_invrnt_dst], $D$ is hyperconvex iff every boundary point of $D$ admits a local psh barrier boundary.
2.8. Growth Vectors and Tangent Cone
In this subsection, we recall the definitions and related results concerning growth vectors and tangent cones from [Whitney].
Definition 2.52.
A set $C\subset\mathbb{C}^{m}$ is an open complex cone if $z\in C$ implies $\lambda z\in C$ for all $\lambda\in\mathbb{C}$.
We may construct open complex cones by taking an arbitrary set $S\subset\mathbb{C}^{N}\setminus\{0\}$ and considering $C=C(S)=\{\lambda z|z\in S,~\lambda\in\Delta\}$. We then refer to $C$ as the cone generated by $S$. The associated punctured (complex) cone is given by $C^{*}=\{\lambda z|z\in S,~\lambda\in\Delta^{*}\}$. Note that if $S$ is convex, then $C(S)$ is a convex cone.
Definition 2.53.
Suppose $f$ is holomorphic near $a$, and ord $_{a}f\geqq v$. Then we may write
(2.1) $f(x)=\sum_{|\lambda|=v}A_{\lambda}\left(x_{1}-a_{1}\right)^{\lambda_{1}}\cdots\left(x_{n}-a_{n}\right)^{\lambda_{n}}+\sum_{|\lambda|=v+1}F_{\lambda}(x)\left(x_{1}-a_{1}\right)^{\lambda_{1}}\cdots\left(x_{n}-a_{n}\right)^{\lambda_{n}}$
where the $F_{\lambda}$ is holomorphic near $a$. The first sum on the right of (2.1) is the initial polynomial of $f$ at $a$; call it $f_{a}^{[*]}(x-a)$.
As a consequence of (2.1), consider any vector $v\in\mathbb{C}^{n}$, use $x=a+zv,z\in\mathbb{C}$ for sufficiently small. Then we have the expansion
$f(a+zv)=z^{v}\left[f_{a}^{[*]}(v)+zR(a,z,v)\right].$
where $R$ is holomorphic for small $z$.
Definition 2.54.
Let $f$ be holomorphic and not identically zero near $a$, with $f(a)=0$. A non-zero vector $v$ is a growth vector for $f$ at $a$ if for the function $g(z):=f(a+zv)$ defined on sufficiently small disk $\Delta_{r}\subset\mathbb{C}$, we have $\operatorname{ord}_{0}g=\operatorname{ord}_{a}f$.
Proposition 2.55.
With $f$ as in the above definition 2.54, let $V_{f}$ be the set of non-growth vectors for $f$ at $a$. Then $V_{f}$ is the set of solutions of $f_{a}^{[*]}(v)=0$; hence it is an algebraic cone, and is closed and nowhere dense in $\mathbb{C}^{N}$.
Remark 2.56.
Note that $dim(V_{f})=2N-2$, because by above proposition, $V_{f}$ is the zero set of homogeneous polynomial $f_{a}^{[*]}$.
Definition 2.57.
Let $V$ be an analytic variety in $\mathbb{C}^{N}$. The tangent cone $C(V,p)$ is the set of all vectors $v$ with the property that for each $\varepsilon>0$ there is a point $q\in V$ and there is a complex number $\alpha\in\mathbb{C}$ such that
$|q-p|<\varepsilon,\quad|\alpha(q-p)-v|<\varepsilon$
Equivalently, $v\in C(V,p)$ if and only if there exist sequences $(q_{i})$ and $(\alpha_{i})$ such that
$q_{i}\in V,\quad q_{i}\rightarrow p,\quad\alpha_{i}\left(q_{i}-p\right)\rightarrow v$
Note that the tangent cone is a complex cone. Thus the direction of (non-zero) $v$ is the limit of directions of secants from $p$ to points of $V$ converging to $p$.
The following theorem establishes the relationship between non-growth vectors of $f$ at $p$ and the tangent cone to the zero set of $f$ at $p$.
Theorem 2.58.
Let $f$ be holomorphic and not identically zero near $p$ with $Z$ as its zero set and $\operatorname{ord}_{p}f=m$. For a non-zero vector $v\in\mathbb{C}^{N}$, define the function
$g_{v}(t)=f(p+tv),~t\in\Delta_{r},$
where $\Delta_{r}$ is a disc of sufficiently small radius $r>0$. Then, $\operatorname{ord}_{0}g_{v}>m$ if and only if $v\in C(Z,p)$, the tangent cone to $Z$ at $p$.
2.9. Analytic and Semi-analytic sets
In this subsection, we shall recall the definitions and results regarding algebraic and analytic sets (varieties) over one of the fields $\mathbb{R}$ or $\mathbb{C}$), from [Singularity] and [Strat]. Throughout this subsection, the term analytic refers to real analytic when $K=\mathbb{R}$, and complex analytic when $K=\mathbb{C}$. Let $M$ denote a connected $n$-dimensional (real or complex) analytic manifold and let $\mathcal{O}(M)$ denote the ring of analytic functions from $M$ to $K$.
Definition 2.59.
Let $A$ be an analytic subset of $M$. A point $p\in A$ is called smooth, of dimension $d$, if and only if there exists an open neighbourhood $W$ of $p$ in $M$ such that $W\cap A$ is an analytic submanifold of $W$ of dimension $d$ (over the field $K$ ). Thus, $p\in A$ is smooth of dimension $d$ if and only if there exists an open neighborhood $W$ of $p$ in $M$ and $f_{d+1},\ldots f_{n}\in\mathcal{O}(W)$ such that $W\cap A=Z\left(f_{d+1},\ldots,f_{n}\right)$ and, for all $x\in W,\nabla f_{d+1}(x),\ldots,\nabla f_{n}(x)$ are linearly independent.
We denote the set of smooth points of $A$ by $\stackrel{{\scriptstyle\circ}}{{A}}$, and the set of smooth points of dimension $d$ by $\stackrel{{\scriptstyle\circ}}{{A}}^{(d)}$. A smooth point of $A$ of highest dimension is called a regular point (this definition of regular is not followed by all authors); we denote the set of regular points of $A$ by $A^{\text{reg }}$. A point $p\in A$ which is not smooth is called a singular point (or, a singularity). We denote the set of singular points, the singular locus, of $A$ by $\Sigma A$.
Theorem 2.60.
Let $A$ be an analytic subset of $M$. Suppose that $0\leq d\leq n$. Then
  1. (1)
    $\stackrel{{\scriptstyle\circ}}{{A}}^{(d)}$ is a d-dimensional analytic submanifold of $M$, and is an open subset of $A$;
  2. (2)
    the $\stackrel{{\scriptstyle\circ}}{{A}}^{(d)}$ are disjoint for different $d,\stackrel{{\scriptstyle\circ}}{{A}}=\stackrel{{\scriptstyle\circ}}{{A}}^{(0)}\cup\ldots\stackrel{{\scriptstyle\circ}}{{A}}^{(n)},\stackrel{{\scriptstyle\circ}}{{A}}$ is an analytic submanifold of $M$, and $\stackrel{{\scriptstyle\circ}}{{A}}$ is open in $X$;
  3. (3)
    $\stackrel{{\scriptstyle\circ}}{{A}}$ is dense in $A$; and
  4. (4)
    $\Sigma A$ is a closed, nowhere dense subset of $A$.
Definition 2.61.
The dimension (over $K$), denoted $\operatorname{dim}A$, of an analytic set $A\subseteq M$ is the largest $d$ such that $\stackrel{{\scriptstyle\circ}}{{A}}^{(d)}$ is non-empty. The codimension of $A$ is the number $\dim M-\dim A$. The dimension of $A$ at a point $p\in A$, denoted $\operatorname{dim}_{p}A$, is the largest $d$ such that $p$ is in the closure of $\stackrel{{\scriptstyle\circ}}{{A}}^{(d)}$.
Definition 2.62.
A subset $V$ of $M$ is called a semivariety if locally at each point $x\in V$ it is a finite union of subsets defined by equations and inequalities
$f_{1}=\cdots=f_{k}=0\quad\begin{cases}g_{1}\neq 0,\ldots,g_{l}\neq 0&\text{ (complex case) },\\ g_{1}>0,\ldots,g_{l}>0&\text{ (real case) },\end{cases}$
where $f_{j}$ ’s and $g_{k}$ ’s are real (or complex) analytic (or algebraic) depending on the case under consideration.
In the real algebraic case, semivarieties are commonly referred to as semialgebraic sets; in the complex algebraic case, they are known as constructible, and in either analytic case, they are called semianalytic sets.
Definition 2.63.
Let $V$ be a semianalytic subset of $M$. We then define smooth points, regular points and singular points exactly as in the above Definition 2.59 for analytic subsets, using the same notations.
In the case of semivarieties, it is known that the smooth points of a semianalytic set are dense. Therefore, we define the dimension of $X$, and the dimension of $X$ at a point exactly as in the above Definition 2.61.
3. Proofs of theorems 1.3, 1.7, 1.15, 1.17
One of our main goals of this section is to characterize the holomorphic retracts of the Cartesian product $B_{1}\times\ldots\times B_{N}$ where the $B_{j}$’s are bounded balanced domains of holomorphy. The result seems to be new at-least to the extent of not being noted elsewhere; in-fact new even in the case when all the $B_{j}$’s are convex in which case the above product will be referred to as a polyball for short. However we hasten to admit at the outset here, that the proof of theorem 1.17 about the aforementioned goal, closely follows the recent work [BBMV-union]; with the new ingredient being theorem 3.6, an analogue of Mazet’s Schwarz lemma for the non-convex case, all discussed below. We give here a thorough treatment despite repetition of some of the essential ideas as it allows for a better exposition of the proof of both this theorem, as well as proofs of theorems prior to it as in the title of this section, which are fundamental to this article.
Recall that theorem 1.17 has $2$ parts, wherein its part (A) requires the boundaries of all the factors of the product domain to satisfy an extremality condition. We discuss some examples for this now. Since the result is new even for the strictly convex case, we shall dwell upon this special case of interest first. Thus until further notice, suppose the $B_{j}$’s are bounded balanced strictly convex domains or equivalently, the unit ball with respect to some norm in $\mathbb{C}^{N_{j}}$ for some $N_{j}\in\mathbb{N}$, for each $j=1,\ldots,N$; as noted in the previous section, this is equivalent to their boundaries being $\mathbb{R}$-extremal and implies (owing to the temporary assumption of convexity) the holomorphic extremality condition required by 1.17 (A). As it turns out that the essential arguments for the case of a product of a pair of such balls extend to the more general case, we shall for simplicity of exposition, focus attention for the most part on the case of a product of $2$ such balls. So, let $B_{1}\subset\mathbb{C}^{N_{1}},B_{2}\subset\mathbb{C}^{N_{2}}$ (where $N_{1},N_{2}\in\mathbb{N}$ with say, $\min\{N_{1},N_{2}\}=N_{1}$) denote a pair of bounded balanced convex domains with $\mathbb{R}$-extremal boundaries; in-particular, $B_{1},B_{2}$ are balls with respect to some norms (on $\mathbb{C}^{N_{1}},\mathbb{C}^{N_{2}}$ respectively) and needless to say are domains of holomorphy. The simplest example of such a domain is the standard unit disc in $\mathbb{C}$ which we denote by $\Delta$. However, many nice properties of the disc may well not hold for such domains (i.e., bounded balanced convex domains). To begin with, balls in the $l^{p}$-norm for $1<p<\infty$, while satisfying the condition that their boundary points are $\mathbb{R}$-extremal, need not always be infinitely smooth – they are atmost $[p]$-smooth, where $\cdot$] denotes the greatest integer function. They are thus atleast $C^{1}$-smooth and to get an example where this $C^{1}$-smoothness fails, we may take an intersection of a pair of suitable dilates of the ball in the $l^{p_{1}}$ norm with that of the $l^{p_{2}}$ norm, for a pair of positive distinct reals $p_{1},p_{2}$ in the interval $(1,\infty)$ to get examples of domains whose boundaries are not $C^{1}$-smooth but nevertheless satisfy the $\mathbb{R}$-extremal condition; as this is essentially the well-known fact that the intersection of a pair of strictly convex domains is again strictly convex, we shall provide details with $\mathbb{R}$-extremality replaced by $\mathbb{C}$-extremality, in the following.
Proposition 3.1.